Find a polynomial equation with integral coefficients that has the roots:
-2, 2, -3

Found 3 solutions by solver91311, MEYDA, MathTherapy:
Answer by solver91311(24713)



A polynomial function has a zero if and only if is a factor of the polynomial.

Hence:



Just multiply it out to get your polynomial.


John


Answer by MEYDA(1)

Hence :
(x-(-2))(x-(-1))(x-3) = 0
(x+2)(x+1)(x-3)= 0
x^3+3x^2-4x-12=0

Answer by MathTherapy(10858)

Find a polynomial equation with integral coefficients that has the roots:
-2, 2, -3
If roots are - 2, 2, and - 3, factors are x + 2, x - 2, and x + 3, and so, we get the CORRECT ANSWER as: (x + 2)(x - 2)(x + 3) = 0. 
Multiplying the binomials to get a 3rd degree equation.